# bibliography.bib

### v3

@@ 1. Introduction
One of the most important problems in representation theory is understanding the category of polynomial representations of the general linear group $\operatorname{GL}(n)$, as well as the limit as $n\to\infty$ of these categories – the category of polynomial functors ([macdonald_symmetric_1995],[friedlander_cohomology_1997]). While over algebraically closed fields of characteristic zero representations of $GL(n)$ are fully understood through the classical work of Schur and Weyl, in positive characteristic the situation is much more complicated, and even computing the dimensions of irreducible representations is very difficult in general. However, the classification of irreducible representations is characteristic free - they are classified by highest weights $(\lambda_{1},...,\lambda_{n})$ where $\lambda_{1}\geq...\geq\lambda_{n}$, $\lambda_{i}\in\mathbb{Z}$ ([jantzen_algebraic_groups_2003]). The story is similar for the general linear supergroups $\operatorname{GL}(m|n)$ and the corresponding categories of superpolynomial functors ([axtell_superpolynomial_2013],[drupieski_cohomology_2016]) - the category of representations of $\operatorname{GL}(m|n)$ is rather complicated already in characteristic zero and much more so in positive characteristic, but the classification of irreducible representations by highest weights is available and uniform in all characteristics ([cheng_wang_superalgebras_2012]).
@@
Recently there has been significant interest in generalizations of this theory to symmetric tensor categories $\mathcal{C}$ of moderate growth, i.e., studying the representation category of the affine group scheme $\operatorname{GL}(X)$ for $X\in\mathcal{C}$ and the corresponding category of polynomial functors obtained by taking an appropriate limit with respect to $X$ (see e.g. [venkatesh_representations_2022],[coulembier_inductive_2024]). It is easy to show that if $F:\mathcal{C}\to\mathcal{D}$ is a tensor functor then the representation category of $\operatorname{GL}(X)$ in $\mathcal{C}$ (compatible with the fundamental group of $\mathcal{C}$) is canonically equivalent to one for $\operatorname{GL}(F(X))$ in $\mathcal{D}$. So, in characteristic zero, since by Deligne’s theorem ([deligne_tensorielles_2002]) there exists $F:\mathcal{C}\to\operatorname{sVec}$, there is nothing new. However, in positive characteristic there are many new possibilities. For example, one can consider this problem in the Verlinde category $\operatorname{Ver}_{p}$. In this case, the classification of irreducible representations of $\operatorname{GL}(X)$ was obtained by Venkatesh [venkatesh_representations_2022]. As the tensor categories admitting a tensor functor $F:\mathcal{C}\to\operatorname{Ver}_{p}$ are precisely the Frobenius exact categories ([coulembier_frobenius_2022]), this takes care of the classification of irreducible representations of $\operatorname{GL}(X)$ (and thereby of simple polynomial functors) in such categories.
@@
However, there are many tensor categories that are not Frobenius exact, for example the categories $\operatorname{Ver}_{p^{n}}^{+}$, $\operatorname{Ver}_{p^{n}}$ defined in [benson_symmetric_2020] in characteristic 2 and in [benson_new_2021],[coulembier_monoidal_2021] in general (namely, $\operatorname{Ver}_{p^{n}}$ is the abelian envelope of the quotient of the category of tilting modules for $\operatorname{SL}_{2}(k)$ by the $n$-th Steinberg module, while $\operatorname{Ver}_{p^{n}}^{+}$ is its subcategory generated by $\operatorname{PGL}_{2}(k)$-modules) when $n\geq 2$. Moreover, it is conjectured [benson_new_2021] that any moderate growth tensor category $\mathcal{C}$ in characteristic $p$ admits a tensor functor $F:\mathcal{C}\to\operatorname{Ver}_{p^{\infty}}:=\operatorname{colim}_{n}\operatorname{Ver}_{p^{n}}$, so it is conjecturally sufficient to understand these examples. The goal of this paper is to do so in the simplest of them, $\mathcal{C}=\operatorname{Ver}_{4}^{+}$, and more generally to develop Lie theory in this category.
@@
Specifically, we develop representation theory of general linear groups in the category $\operatorname{Ver}_{4}^{+}$, the tensor category of modules over the Hopf algebra $k[d]/d^{2}$ with $d$ primitive and $\text{char}(k)=2$, with commutativity defined by the $R$-matrix $R:=1\otimes 1+d\otimes d$ ([venkatesh_hilbert_2016]). [^1] Note that this category has implicitly appeared much earlier than the referenced papers on Verlinde categories in the context of homotopy theory, see [etingof_p-adic_2020], Remark 3.4. Since $\operatorname{Ver}_{4}^{+}$ is a reduction of the category of supervector spaces to characteristic $2$ ([venkatesh_hilbert_2016], Subsection 2.3), these groups may be viewed as general linear supergroups in characteristic $2$, and more generally Lie theory in $\operatorname{Ver}_{4}^{+}$ may be regarded as a nontrivial version of superLie theory in characteristic $2$ (the trivial one being ordinary Lie theory, as there are no signs). More precisely, every object in $\operatorname{Ver}_{4}^{+}$ has the form $m\mathds{1}+nP$ where $P$ is the indecomposable projective, and $\operatorname{GL}(m\mathds{1}+nP)$ is the reduction to characteristic $2$ of $\operatorname{GL}(m+n|n)$. We explicitly describe the irreducible representations of $\operatorname{GL}(P)$ and then use this description to classify the irreducible representations of $\operatorname{GL}(m\mathds{1}+nP)$ for general $m,n$. We also define some subgroups of $\operatorname{GL}(m\mathds{1}+nP)$ and classify their irreducible representations. Finally, we conjecture a Steinberg tensor product theorem for $\operatorname{Ver}_{4}^{+}$ involving the square of the Frobenius map.
@@
We note that there are other nontrivial versions of superLie theory in characteristic $2$, for example the one discussed in [bouarroudj_vectorial_2020] (see Subsection 1.2.3) and [bouarroudj_classification_2023] (see Subsection 2.2). In this version, Lie superalgebras are defined as $\mathbb{Z}/2\mathbb{Z}$-graded Lie algebras, with an additional structure (the squaring map), whereas in our setting Lie superalgebras have no $\mathbb{Z}/2\mathbb{Z}$-grading, and instead have a derivation $d$ such that $d^{2}=0$.
@@ 2.1. Symmetric Tensor Categories
We refer to [etingof_tensor_2015] for a full introduction to the theory of tensor categories.
@@
(Symmetric) tensor categories are a natural generalization of representation categories for affine group and supergroup schemes. For a more detailed discussion of symmetric tensor categories, we refer the reader to [etingof_lectures_2021].
@@
We refer to [etingof_frobenius_2020] and [coulembier_frobenius_2022] for a discussion on Frobenius functors in general symmetric tensor categories. In characteristic 2, following [etingof_frobenius_2020], we can define the Frobenius functor as
@@
If $\operatorname{Fr}$ is exact, we say that $\mathcal{C}$ is Frobenius exact. In [coulembier_frobenius_2022], it is proven that a Frobenius exact symmetric tensor category of moderate growth over a field of characteristic 2 is equivalent to the representation category of some affine group scheme.
@@ 2.2. Commutative algebras and (affine) group schemes in tensor categories
An associative algebra $(A,\mu,\eta)$ in a symmetric tensor category $\mathcal{C}$ is an object $A\in\operatorname{ind}\mathcal{C}$ equipped with a multiplication map $\mu:A\otimes A\to A$ and a unit map $\eta:\mathds{1}\to A$ satisfying the standard associativity and unit axioms. A commutative algebra is an associative algebra that also satisfies the commutativity axioms for $\mu$.  [^2] See [coulembier_commutative_2023] for an overview of commutative algebra in STCs and [coulembier_algebraic_2023] for an overview of algebraic geometry in symmetric tensor categories. We denote the category of commutative algebras in $\mathcal{C}$ by $\operatorname{CommAlg}\mathcal{C}$.
@@ 2.2.1. The fundamental group of a symmetric tensor category
Recall that every symmetric tensor category $\mathcal{C}$ has a canonical affine group scheme associated to it, its fundamental group $\pi_{1}(\mathcal{C})$. As a functor of points, $\pi_{1}(\mathcal{C})$ is defined by $\pi_{1}(\mathcal{C})(A)=\operatorname{Aut}^{\otimes}(A\otimes-)$ (that is, tensor automorphisms of the functor $A\otimes-$). The fundamental group was introduced in [deligne_categories_1990]. $\pi_{1}(\mathcal{C})$ has a natural action on every object $X\in\mathcal{C}$ since $\pi_{1}(\mathcal{C})(A)$ acts on $A\otimes X$.
@@ 2.2.2. Distribution Algebras
The proof that $\operatorname{Dist}(G)$ is a cocommutative Hopf algebra is identical to the proof in [venkatesh_harish-chandra_2022] Lemma 4.31. $(I/I^{2})^{*}$ is a Lie subalgebra of $\operatorname{Dist}(G)$ because it is the maximal primitive subobject of $\operatorname{Dist}(G)$, as primitivity is equivalent to vanishing on $1$ and on $I^{2}$. ∎
@@ 2.3. The category $\operatorname{Ver}_{4}^{+}$
Let $k$ be a field of characteristic 2, not necessarily algebraically closed. The categories $\operatorname{Ver}_{2^{n}}(k),\operatorname{Ver}_{2^{n}}^{+}(k)$ were introduced in [benson_symmetric_2020]. For $n=2$, $\operatorname{Ver}_{4}^{+}(k)$ was known before as the reduction to characteristic $2$ of $\operatorname{sVec}$ ([venkatesh_hilbert_2016]). Also, $\operatorname{Ver}_{4}^{+}(k)$ has a concrete description. Consider the Hopf algebra $k[d]/d^{2}$ where $d$ is primitive; then $\operatorname{Ver}_{4}^{+}(k)$ is the monoidal category of finite-dimensional $k[d]/d^{2}$-modules with symmetric braiding given by $1\otimes 1+d\otimes d$, i.e.
@@ 2.3.1. $\operatorname{Ver}_{4}^{+}$ as reduction of $\operatorname{sVec}$ to characteristic 2
In this section, let $K:=\mathbb{Q}_{2}(\sqrt{2})$, where $\mathbb{Q}_{2}$ is the field of $2$-adic numbers. We recall how $\operatorname{Ver}_{4}^{+}(\mathbb{F}_{2})$ can be constructed as a non-semisimple reduction of $\operatorname{sVec}(K)$ over $K$. This construction was described in [venkatesh_hilbert_2016], Section 2.3. Let $H:=K[\mathbb{Z}/2\mathbb{Z}]$ be the group algebra of $\mathbb{Z}/2\mathbb{Z}$ as a Hopf algebra, i.e. it is generated by $1,g$ such that $g^{2}=1$ and $\Delta(g)=g\otimes g$. Recall that in characteristic not 2, the category of supervector spaces is the $H$-modules $\operatorname{Rep}(H)$ with associated $R$-matrix
@@
If $v$ is even, $gv=v$, and if $w$ is odd, $gw=-w$. Since $d=\frac{1-g}{\sqrt{2}}$, we have $dv=0$ and $dw=\sqrt{2}w$. Therefore, for $V=k^{(1,0)}=\langle v\rangle$, its reduction modulo $\sqrt{2}$ is $\mathds{1}$.
@@ 2.3.2. The fundamental group of $\operatorname{Ver}_{4}^{+}$
Suppose we have an isomorphism $P\otimes P\cong 2P$ sending $1\otimes 1\mapsto x$, $d\otimes 1\mapsto y$, so $1\otimes d+d\otimes 1\mapsto x^{\prime}$ and $d\otimes d\mapsto y^{\prime}$. Since $\eta_{P\otimes P}$ must fix elements in $\operatorname{im}d$, in particular $1\otimes d+d\otimes 1$, we must have $11+b^{\prime}=$ 1, so $b^{\prime}=0$. Also, we want
@@
and likewise for $1\otimes d$. Therefore, $\eta_{P}=\begin{pmatrix}1&0\\ b&1\end{pmatrix}$ where $b^{2}=0$, $b^{\prime}=0$, so $\eta$ is parametrized by elements in $\ker d$ that square to 00.
@@ 2.4. Highest Weights
From this section onwards, we assume our symmetric tensor category $\mathcal{C}$ satisfies hypothesis 1.3.1 in [coulembier_algebraic_2023], which allows us to develop scheme theory in $\mathcal{C}$. In particular, $\operatorname{Ver}_{4}^{+}$ satisfies this hypothesis, see [coulembier_algebraic_2023] Section 8.
@@ 3.2. $\mathbb{G}_{a},\mathbb{G}_{m}$, and friends
The coordinate ring of $\mathbb{G}_{m}$ is $k[T,T^{-1}]$, $T$ grouplike. The coordinate ring of $\mathbb{G}^{\prime}_{m}$ is $k[T,T^{-1},T^{\prime}]$, $T$ grouplike. Its Lie algebra’s underlying object is again $P$, but it has a nonzero bracket, $[x,x]=x^{\prime}$.
@@
$M_{1}$ has coordinate ring $k[X,X^{\prime}]$, $\Delta(X)=X\otimes 1+1\otimes X+X^{\prime}\otimes X$, $S(X)=X+XX^{\prime}$. Therefore, $1+X^{\prime}$ is grouplike. The Lie algebra of this group has underlying object $P$, with nonzero brackets $[x,x^{\prime}]=[x^{\prime},x]=x^{\prime}$.
Since $\operatorname{Ver}_{4}^{+}$ is a reduction of $\operatorname{sVec}$ to characteristic 2, let us compare this to the corresponding groups in characteristic 00 in $\operatorname{sVec}$. For the rest of the subsection, we let $k=\mathbb{Q}_{2}(\sqrt{2})$. The following two propositions are well-known statements about low-dimensional Lie superalgebras and supergroups.
@@
The Lie algebra structure on $k^{1|1}$ with $x$ even and $\xi$ odd and nonzero bracket $[x,\xi]=\xi$ reduces to the Lie algebra of $M_{1}$. Consider $v=x+\xi$ and $w=\sqrt{2}\xi$. Then $dv=w$ and $dw=\sqrt{2}w$. Moreover, $[v,v]=0$, $[v,w]=w$, and $[w,w]=0$. Reducing modulo $\sqrt{2}$, we have $dv=w$, $dw=0$, and $[v,w]=w$ is the only nonzero bracket, which is $\mathfrak{aff}(0,1)$, the Lie algebra of the group $\operatorname{Aff}(0,1)$ of affine transformations of $k^{(0,1)}$.
@@
The $(1|1)$-dimensional supersymmetry group $S(1|1)$ taking a supercommutative algebra $A$ to $A^{\times}$ in characteristic 00 degenerates to $\mathbb{G}^{\prime}_{m}$ in characteristic 2.
@@
Let $k=\mathbb{Q}_{2}[\sqrt{2}]$ and $k[\mathbb{Z}/2\mathbb{Z}]=\langle 1,g\rangle$. Let $A$ be a supercommutative algebra. Every invertible element of $A$ can be written uniquely as $a+\alpha$, $a$ even and invertible, $\alpha$ odd (such an element has inverse $a-\frac{\alpha}{a^{2}}$). Thus, $\mathcal{O}(G)=k[X,Y,X^{-1}]$, where $X$ is even, $Y$ is odd, and
@@
Since $X$ is even and $Y$ is odd, $gX=X,gY=-Y$. Moreover, $d=\frac{1-g}{\sqrt{2}}$, so $dX=0$ and $dY=\sqrt{2}Y$. Let $p=X+Y$ and $q=\sqrt{2}Y$, so
@@
The Lie superalgebra of this group has only one nonzero bracket, $[\mu_{Y},\mu_{Y}]$, where $\mu_{Y}$ is the dual element to $Y$. This Lie superalgebra is $\mathfrak{s}(1|1)$, the Lie algebra of $S(1|1)$.

### v4

# Representation Theory of General Linear Supergroups in Characteristic 2

@@ 1. Introduction
One of the most important problems in representation theory is understanding the category of polynomial representations of the general linear group $\operatorname{GL}(n)$, as well as the limit as $n\to\infty$ of these categories – the category of polynomial functors ([23],[19]). While over algebraically closed fields of characteristic zero representations of $GL(n)$ are fully understood through the classical work of Schur and Weyl, in positive characteristic the situation is much more complicated, and even computing the dimensions of irreducible representations is very difficult in general. However, the classification of irreducible representations is characteristic free - they are classified by highest weights $(\lambda_{1},...,\lambda_{n})$ where $\lambda_{1}\geq...\geq\lambda_{n}$, $\lambda_{i}\in\mathbb{Z}$ ([20]). The story is similar for the general linear supergroups $\operatorname{GL}(m|n)$ and the corresponding categories of superpolynomial functors ([1],[14]) - the category of representations of $\operatorname{GL}(m|n)$ is rather complicated already in characteristic zero and much more so in positive characteristic, but the classification of irreducible representations by highest weights is available and uniform in all characteristics ([6]).
@@
Recently there has been significant interest in generalizations of this theory to symmetric tensor categories $\mathcal{C}$ of moderate growth, i.e., studying the representation category of the affine group scheme $\operatorname{GL}(X)$ for $X\in\mathcal{C}$ and the corresponding category of polynomial functors obtained by taking an appropriate limit with respect to $X$ (see e.g. [26],[9]). It is easy to show that if $F:\mathcal{C}\to\mathcal{D}$ is a tensor functor then the representation category of $\operatorname{GL}(X)$ in $\mathcal{C}$ (compatible with the fundamental group of $\mathcal{C}$) is canonically equivalent to one for $\operatorname{GL}(F(X))$ in $\mathcal{D}$. So, in characteristic zero, since by Deligne’s theorem ([12]) there exists $F:\mathcal{C}\to\operatorname{sVec}$, there is nothing new. However, in positive characteristic there are many new possibilities. For example, one can consider this problem in the Verlinde category $\operatorname{Ver}_{p}$. In this case, the classification of irreducible representations of $\operatorname{GL}(X)$ was obtained by Venkatesh [26]. As the tensor categories admitting a tensor functor $F:\mathcal{C}\to\operatorname{Ver}_{p}$ are precisely the Frobenius exact categories ([11]), this takes care of the classification of irreducible representations of $\operatorname{GL}(X)$ (and thereby of simple polynomial functors) in such categories.
@@
However, there are many tensor categories that are not Frobenius exact, for example the categories $\operatorname{Ver}_{p^{n}}^{+}$, $\operatorname{Ver}_{p^{n}}$ defined in [2] in characteristic 2 and in [3],[10] in general (namely, $\operatorname{Ver}_{p^{n}}$ is the abelian envelope of the quotient of the category of tilting modules for $\operatorname{SL}_{2}(k)$ by the $n$-th Steinberg module, while $\operatorname{Ver}_{p^{n}}^{+}$ is its subcategory generated by $\operatorname{PGL}_{2}(k)$-modules) when $n\geq 2$. Moreover, it is conjectured [3] that any moderate growth tensor category $\mathcal{C}$ in characteristic $p$ admits a tensor functor $F:\mathcal{C}\to\operatorname{Ver}_{p^{\infty}}:=\operatorname{colim}_{n}\operatorname{Ver}_{p^{n}}$, so it is conjecturally sufficient to understand these examples. The goal of this paper is to do so in the simplest of them, $\mathcal{C}=\operatorname{Ver}_{4}^{+}$, and more generally to develop Lie theory in this category.
@@
Specifically, we develop representation theory of general linear groups in the category $\operatorname{Ver}_{4}^{+}$, the tensor category of modules over the Hopf algebra $k[d]/d^{2}$ with $d$ primitive and $\text{char}(k)=2$, with commutativity defined by the $R$-matrix $R:=1\otimes 1+d\otimes d$ ([25]). [^1] Note that this category has implicitly appeared much earlier than the referenced papers on Verlinde categories in the context of homotopy theory, see [16], Remark 3.4. Since $\operatorname{Ver}_{4}^{+}$ is a reduction of the category of supervector spaces to characteristic $2$ ([25], Subsection 2.3), these groups may be viewed as general linear supergroups in characteristic $2$, and more generally Lie theory in $\operatorname{Ver}_{4}^{+}$ may be regarded as a nontrivial version of superLie theory in characteristic $2$ (the trivial one being ordinary Lie theory, as there are no signs). More precisely, every object in $\operatorname{Ver}_{4}^{+}$ has the form $m\mathds{1}+nP$ where $P$ is the indecomposable projective, and $\operatorname{GL}(m\mathds{1}+nP)$ is the reduction to characteristic $2$ of $\operatorname{GL}(m+n|n)$. We explicitly describe the irreducible representations of $\operatorname{GL}(P)$ and then use this description to classify the irreducible representations of $\operatorname{GL}(m\mathds{1}+nP)$ for general $m,n$. We also define some subgroups of $\operatorname{GL}(m\mathds{1}+nP)$ and classify their irreducible representations. Finally, we conjecture a Steinberg tensor product theorem for $\operatorname{Ver}_{4}^{+}$ involving the square of the Frobenius map.
@@
We note that there are other nontrivial versions of superLie theory in characteristic $2$, for example the one discussed in [4] (see Subsection 1.2.3) and [5] (see Subsection 2.2). In this version, Lie superalgebras are defined as $\mathbb{Z}/2\mathbb{Z}$-graded Lie algebras, with an additional structure (the squaring map), whereas in our setting Lie superalgebras have no $\mathbb{Z}/2\mathbb{Z}$-grading, and instead have a derivation $d$ such that $d^{2}=0$.
@@ 2.1. Symmetric Tensor Categories
We refer to [15] for a full introduction to the theory of tensor categories.
@@
(Symmetric) tensor categories are a natural generalization of representation categories for affine group and supergroup schemes. For a more detailed discussion of symmetric tensor categories, we refer the reader to [17].
@@
We refer to [18] and [11] for a discussion on Frobenius functors in general symmetric tensor categories. In characteristic 2, following [18], we can define the Frobenius functor as
@@
If $\operatorname{Fr}$ is exact, we say that $\mathcal{C}$ is Frobenius exact. In [11], it is proven that a Frobenius exact symmetric tensor category of moderate growth over a field of characteristic 2 is equivalent to the representation category of some affine group scheme.
@@ 2.2. Commutative algebras and (affine) group schemes in tensor categories
An associative algebra $(A,\mu,\eta)$ in a symmetric tensor category $\mathcal{C}$ is an object $A\in\operatorname{ind}\mathcal{C}$ equipped with a multiplication map $\mu:A\otimes A\to A$ and a unit map $\eta:\mathds{1}\to A$ satisfying the standard associativity and unit axioms. A commutative algebra is an associative algebra that also satisfies the commutativity axioms for $\mu$.  [^2] See [8] for an overview of commutative algebra in STCs and [7] for an overview of algebraic geometry in symmetric tensor categories. We denote the category of commutative algebras in $\mathcal{C}$ by $\operatorname{CommAlg}\mathcal{C}$.
@@ 2.2.1. The fundamental group of a symmetric tensor category
Recall that every symmetric tensor category $\mathcal{C}$ has a canonical affine group scheme associated to it, its fundamental group $\pi_{1}(\mathcal{C})$. As a functor of points, $\pi_{1}(\mathcal{C})$ is defined by $\pi_{1}(\mathcal{C})(A)=\operatorname{Aut}^{\otimes}(A\otimes-)$ (that is, tensor automorphisms of the functor $A\otimes$-). The fundamental group was introduced in [13]. $\pi_{1}(\mathcal{C})$ has a natural action on every object $X\in\mathcal{C}$ since $\pi_{1}(\mathcal{C})(A)$ acts on $A\otimes X$.
@@ 2.2.2. Distribution Algebras
The proof that $\operatorname{Dist}(G)$ is a cocommutative Hopf algebra is identical to the proof in [24] Lemma 4.31. $(I/I^{2})^{*}$ is a Lie subalgebra of $\operatorname{Dist}(G)$ because it is the maximal primitive subobject of $\operatorname{Dist}(G)$, as primitivity is equivalent to vanishing on $1$ and on $I^{2}$. ∎
@@ 2.3. The category $\operatorname{Ver}_{4}^{+}$
Let $k$ be a field of characteristic 2, not necessarily algebraically closed. The categories $\operatorname{Ver}_{2^{n}}(k),\operatorname{Ver}_{2^{n}}^{+}(k)$ were introduced in [2]. For $n=2$, $\operatorname{Ver}_{4}^{+}(k)$ was known before as the reduction to characteristic $2$ of $\operatorname{sVec}$ ([25]). Also, $\operatorname{Ver}_{4}^{+}(k)$ has a concrete description. Consider the Hopf algebra $k[d]/d^{2}$ where $d$ is primitive; then $\operatorname{Ver}_{4}^{+}(k)$ is the monoidal category of finite-dimensional $k[d]/d^{2}$-modules with symmetric braiding given by $1\otimes 1+d\otimes d$, i.e.
@@ 2.3.1. $\operatorname{Ver}_{4}^{+}$ as reduction of $\operatorname{sVec}$ to characteristic 2
In this section, let $K:=\mathbb{Q}_{2}(\sqrt{2})$, where $\mathbb{Q}_{2}$ is the field of $2$-adic numbers. We recall how $\operatorname{Ver}_{4}^{+}(\mathbb{F}_{2})$ can be constructed as a non-semisimple reduction of $\operatorname{sVec}(K)$ over $K$. This construction was described in [25], Section 2.3. Let $H:=K[\mathbb{Z}/2\mathbb{Z}]$ be the group algebra of $\mathbb{Z}/2\mathbb{Z}$ as a Hopf algebra, i.e. it is generated by $1,g$ such that $g^{2}=1$ and $\Delta(g)=g\otimes g$. Recall that in characteristic not 2, the category of supervector spaces is the $H$-modules $\operatorname{Rep}(H)$ with associated $R$-matrix
@@
If $v$ is even, $gv=v$, and if $w$ is odd, gw=wgw=-w. Since $d=\frac{1-g}{\sqrt{2}}$, we have $dv=0$ and $dw=\sqrt{2}w$. Therefore, for $V=k^{(1,0)}=\langle v\rangle$, its reduction modulo $\sqrt{2}$ is $\mathds{1}$.
@@ 2.3.2. The fundamental group of $\operatorname{Ver}_{4}^{+}$
Suppose we have an isomorphism $P\otimes P\cong 2P$ sending $1\otimes 1\mapsto x$, $d\otimes 1\mapsto y$, so $1\otimes d+d\otimes 1\mapsto x^{\prime}$ and $d\otimes d\mapsto y^{\prime}$. Since $\eta_{P\otimes P}$ must fix elements in $\operatorname{im}d$, in particular $1\otimes d+d\otimes 1$, we must have $1+b^{\prime}=1$, so $b^{\prime}=0$. Also, we want
@@
and likewise for $1\otimes d$. Therefore, $\eta_{P}=\begin{pmatrix}1&0\\ b&1\end{pmatrix}$ where $b^{2}=0$, $b^{\prime}=0$, so $\eta$ is parametrized by elements in $\ker d$ that square to 0.
@@ 2.4. Highest Weights
From this section onwards, we assume our symmetric tensor category $\mathcal{C}$ satisfies hypothesis 1.3.1 in [7], which allows us to develop scheme theory in $\mathcal{C}$. In particular, $\operatorname{Ver}_{4}^{+}$ satisfies this hypothesis, see [7] Section 8.
@@ 3.2. $\mathbb{G}_{a},\mathbb{G}_{m}$, and friends
The coordinate ring of $\mathbb{G}_{m}$ is $k[T,T^{-1}]$, $T$ grouplike. The coordinate ring of $\mathbb{G}^{\prime}_{m}$ is $k[T,T^{-1},T^{\prime}]$, $T$ grouplike. Its Lie algebra’s underlying object is again $P$, but it has a nonzero bracket, $x,x]=x^{\prime}$.
@@
$M_{1}$ has coordinate ring $k[X,X^{\prime}]$, $\Delta(X)=X\otimes 1+1\otimes X+X^{\prime}\otimes X$, $S(X)=X+XX^{\prime}$. Therefore, $1+X^{\prime}$ is grouplike. The Lie algebra of this group has underlying object $P$, with nonzero brackets $x,x^{\prime}]=[x^{\prime},x]=x^{\prime}$.
Since $\operatorname{Ver}_{4}^{+}$ is a reduction of $\operatorname{sVec}$ to characteristic 2, let us compare this to the corresponding groups in characteristic 0 in $\operatorname{sVec}$. For the rest of the subsection, we let $k=\mathbb{Q}_{2}(\sqrt{2})$. The following two propositions are well-known statements about low-dimensional Lie superalgebras and supergroups.
@@
The Lie algebra structure on $k^{1|1}$ with $x$ even and $\xi$ odd and nonzero bracket $x,\xi]=\xi$ reduces to the Lie algebra of $M_{1}$. Consider $v=x+\xi$ and $w=\sqrt{2}\xi$. Then $dv=w$ and $dw=\sqrt{2}w$. Moreover, $[v,v]=0$, $[v,w]=w$, and $[w,w]=0$. Reducing modulo $\sqrt{2}$, we have $dv=w$, $dw=0$, and $[v,w]=w$ is the only nonzero bracket, which is $\mathfrak{aff}(0,1)$, the Lie algebra of the group $\operatorname{Aff}(0,1)$ of affine transformations of $k^{(0,1)}$.
@@
The $(1|1)$-dimensional supersymmetry group $S(1|1)$ taking a supercommutative algebra $A$ to $A^{\times}$ in characteristic 0 degenerates to $\mathbb{G}^{\prime}_{m}$ in characteristic 2.
@@
Let $k=\mathbb{Q}_{2}[\sqrt{2}]$ and $k[\mathbb{Z}/2\mathbb{Z}]=\langle 1,g\rangle$. Let $A$ be a supercommutative algebra. Every invertible element of $A$ can be written uniquely as $\alpha, aa$ even and invertible, $\alpha$ odd (such an element has inverse $a-\frac{\alpha}{a^{2}}$). Thus, $\mathcal{O}(G)=k[X,Y,X^{-1}]$, where $X$ is even, $Y$ is odd, and
@@
Since $X$ is even and $Y$ is odd, gX=X,gY=YgX=X,gY=-Y. Moreover, $d=\frac{1-g}{\sqrt{2}}$, so $dX=0$ and $dY=\sqrt{2}Y$. Let $p=X+Y$ and $q=\sqrt{2}Y$, so
@@
The Lie superalgebra of this group has only one nonzero bracket, $\mu_{Y},\mu_{Y}]$, where $\mu_{Y}$ is the dual element to $Y$. This Lie superalgebra is $\mathfrak{s}(1|1)$, the Lie algebra of $S(1|1)$.